OpenStudy (anonymous):

At maximum speed, an airplane travels 1720 miles against the wind in 5 hours. Flying with the wind, the plane can travel the same distance in 4 hours. Let x be the maximum speed of the plane and y be the speed of the wind. What is the speed of the plane with no wind?

OpenStudy (anonymous):

1720/5 = Vplane - Vwind (1) 1720/4 = Vplane + Vwind (2)

OpenStudy (anonymous):

344 = x - y 430 = x + y add equations 774 = 2x x = 387mph

OpenStudy (anonymous):

1720 = 5(x - y) 1720 = 4(x + y)

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

x - y = 344 x + y = 430 2x = 774 x = 387 y = 430 - 387 = 43 Maximum speed of aircraft is 387 mph, wind speed is 43 mph.